OpenStudy (anonymous):

Last year, a baseball team paid $20 per bat and $12 per glove, spending a total of $1040. This year, the prices went up to $25 per bat and $16 per glove. The team spent $1350 to purchase the same amount of eqipment as last year. How many bats and gloves did the team purchase each year?

OpenStudy (anonymous):

\[20x+12y=1040\] \[25+16y=1350\]

OpenStudy (anonymous):

\[20x+12y=1040\] \[25x+16y=1350\]

OpenStudy (anonymous):

oh ota, that explains why i kept getting a weird answe. thanks that's all i needed mister ^^